# Fundamental measure theory for predicting many-body correlation functions

### v1

@@ Abstract
We study many-body correlation functions within various Fundamental Measure Theory (FMT) formulations and compare their predictions to Monte Carlo simulations of hard-sphere fluids. FMT accurately captures key features of three- and four-body structure, particularly at low and intermediate wavevectors. We show that the dominant contributions to the four-point structure factor arise from direct triplet correlations, allowing evaluation of four-point correlations to be greatly simplified. In glass-forming liquids at high volume fractions, FMT correctly reproduces deviations from the convolution approximation, highlighting FMT’s ability to capture growing structural multipoint correlations upon supercooling.
@@ Introduction
Here $\alpha$ is a label that identifies the particular weight function. The $n$-body direct correlation functions $c_{n}$ are key quantities in density functional theory and FMT, as they describe the response of the system to density perturbations. These functions are derived from the excess free energy functional. In turn, they are related to the $n$-body density through the Ornstein-Zernike relations.
@@
In the homogeneous single-component bulk phase, the two-, three-, and four-body Ornstein-Zernike relations between $c_{n}$ and $S_{n}$ are
@@
$\begin{split}S_{4}&(\mathbf{k}_{1},\mathbf{k}_{2},\mathbf{k}_{3})=\\ &\left(1+\rho^{2}c_{3}(\mathbf{k}_{1},\mathbf{k}_{2})\right)S_{2}(k_{1})S_{2}(k_{2})S_{3}(\mathbf{k}_{1}+\mathbf{k}_{2},\mathbf{k}_{3})\\ +&\left(1+\rho^{2}c_{3}(\mathbf{k}_{1},\mathbf{k}_{3})\right)S_{2}(k_{1})S_{3}(\mathbf{k}_{1}+\mathbf{k}_{3},\mathbf{k}_{2})S_{2}(k_{3})\\ +&\left(1+\rho^{2}c_{3}(\mathbf{k}_{2},\mathbf{k}_{3})\right)S_{3}(\mathbf{k}_{2}+\mathbf{k}_{3},\mathbf{k}_{1})S_{2}(k_{2})S_{2}(k_{3})\\ -&\left(2-\rho^{3}c_{4}(\mathbf{k}_{1},\mathbf{k}_{2},\mathbf{k}_{3})\right)\\ &\quad\times S_{2}(k_{1})S_{2}(|\mathbf{k}_{1}+\mathbf{k}_{2}+\mathbf{k}_{3}|)S_{2}(k_{2})S_{2}(k_{3}).\end{split}$ (10)
These expressions provide the basis for computing three- and four-body correlations and structure factors with FMT. The higher order structure factors are often approximated purely in terms of $S_{2}$ by assuming that $c_{n>2}=0$, an approach known as the convolution approximation [42, 43, 12].
@@
Kierlik and Rosinberg [44] proposed an alternative formulation that uses only scalar weight functions. In this version, the vector contributions are absorbed into redefined scalar weighted densities so that the same bulk thermodynamics is recovered. Because it recovers exactly the correlation functions generated by Rosenfeld’s original FMT, the formulation of Kierlik and Rosinberg is not treated separately here.
@@
$\begin{split}&\Phi_{\text{T}}=-n_{0}\ln(1-n_{3})+\frac{n_{1}n_{2}-\mathbf{n}_{\mathrm{v}1}\cdot\mathbf{n}_{\mathrm{v}2}}{1-n_{3}}\\ &+\frac{n_{2}^{3}-3n_{2}\,\mathbf{n}_{\mathrm{v}2}\cdot\mathbf{n}_{\mathrm{v}2}+9\left[\mathbf{n}^{T}_{\mathrm{v}2}\mathbf{n}_{\mathrm{m}2}\mathbf{n}_{\mathrm{v}2}-\mathrm{tr}\left({\mathbf{n}_{\mathrm{m}2}^{3}}\right)/2\right]}{24\pi(1-n_{3})^{2}}.\end{split}$ (16)
@@
Later, Malijevskỳ [22] (M) proposed a free energy density that reproduces the equation of state of Boubl´ık [49], showing very accurate density profiles near a hard wall. This free energy density reads
@@
where $j_{0}(x)=\sin(x)/x$, $j_{1}(x)=(\sin(x)-x\cos(x))/x^{2}$, and $j_{2}(x)=(3\sin(x)-3x\cos(x)-x^{2}\sin(x))/x^{3}$ are the first three spherical Bessel functions of the first kind, $\mathbf{r}\mathbf{r}$ and $\mathbf{k}\mathbf{k}$ are dyadic products, and $\mathbf{I}$ is the unit tensor. The Fourier representations of the weight functions are used to construct the direct correlation functions in Fourier space using Eq. (6).
@@ Results
To assess the accuracy of various FMT formulations in capturing $n$-body structural correlations, we compute the direct correlation functions $c_{n}(k)$ and static structure factors $S_{n}(k)$ for a hard-sphere fluid in which the particles have diameter $\sigma=1$, which serves as our unit of length. We consider a number density $\rho=0.94$ just below freezing, corresponding to a volume fraction $\eta=\pi\rho/6\approx 0.49$. For reference, we compare each of the theories to correlation functions measured from standard Metropolis MC simulations of $N=30,000$ particles. We start with a comparison for $n=2$, and consider n=3, 4n=3,\ 4 later in this section.
To place the performance of the different FMT formulations for the 2-point bulk structure in theoretical context, we compare the FMT predictions here additionally to the very successful Verlet (V) [50, 51] and Modified Hypernetted Chain (MHNC) closures of the Ornstein-Zernike relation [52], which are the go-to methods for predicting pair structure in bulk. The MHNC closure uses a parametrization of the bridge function of hard spheres measured from Monte Carlo (MC) simulations and is therefore expected to offer virtually exact predictions [53, 54]. The integral equation theories are solved by the software introduced in Ref. [54].
@@
Figure 1: (a) Fourier-space two-body direct correlation function $\rho c_{2}(k)$ for a hard-sphere fluid at volume fraction $\eta=0.49$, comparing various theoretical models to Monte Carlo (MC) simulation data (black symbols). (b) Magnification of the low-$k$ region. FMT-based predictions are shown for the original Rosenfeld (R), White Bear I (WBI), White Bear II (WBII), and Malijevský (M) functionals, alongside results obtained from Ornstein–Zernike closures, specifically the Verlet (V) and Modified Hypernetted Chain (MHNC) theories.
Figure 1 shows the two‐body direct correlation function $\rho c_{2}(k)$ for a hard‐sphere fluid at volume fraction $\eta=0.49$ comparing Monte Carlo data to a range of theoretical models: Rosenfeld’s original fundamental‐measure theory (R), White Bear I (WBI), White Bear II (WBII), Malijevský’s variant (M), and the two integral‐equation closures. The tensorial theories (T, WBIIt) are not considered here because their tensorial contributions have no effect at the pair level, reducing to their non-tensorial equivalent (R, WBII). Qualitatively, all FMTs show the correct behavior for the direct correlation function. Quantitatively, for the longest wavelengths ($k\to 0$), Rosenfeld’s FMT predicts a considerably more negative $c_{2}(k)$ than the others. This highlights its commonality with Percus–Yevick theory, resulting in an isothermal compressibility that is lower than that computed from MC data [1]. In contrast, the White Bear I/II functionals, which enforce the Boublík–Mansoori–Carnahan–Starling–Leland or generalized Carnahan–Starling equation of state [45, 46, 47], bring $c_{2}(0)$ into near‐perfect agreement with simulation. Malijevský’s free energy density, also built on a high‐accuracy bulk equation of state, gives essentially the same low‐$k$ result as WBII, very slightly underestimating $c_{2}(0)$ compared to WBI/II and the MC data. The Verlet closure, likewise, matches the MC compressibility almost exactly, whereas the MHNC closure fails to accurately capture the small-$k$ behavior because its parametrization was performed for $r>1$ only. At high $k$ ($k\gtrsim 6$), every model’s $c_{2}(k)$ approaches zero in lockstep with MC. To visualize the differences between the theories, we consider the static structure factor $S(k)=1/(1-\rho c_{2}(k))$, which amplifies the small differences in $\rho c(k)$ at larger $k$.
@@
Figure 2 presents the static structure factor $S_{2}(k)$. The principal diffraction peak at $k\approx 7$ attains $S_{\text{max}}\approx 2.9$ in simulation. Rosenfeld overestimates this peak ($S_{\text{max}}\approx 3.1$), while WBI and WBII underestimate it ($S_{\text{max}}\approx 2.7$ and $S_{\text{max}}\approx 2.4$, respectively). Malijevský’s form produces much better agreement with simulation data. The Verlet closure matches the first peak with virtually no error, and so does MHNC.
Figure 3: Triplet structure factor $S_{3}(k_{1},k_{2},\theta)$ at $\eta=0.49$ from Monte Carlo (MC) simulations (black circles) compared to predictions from the different FMTs. The convolution approximation ($c_{3}=0$) is also shown, for which we have used the Rosenfeld prediction of $S_{2}$. Top row: $S_{3}$ as a function of $k_{1}$ at fixed $\theta$ and $k_{2}$. Bottom row: $S_{3}$ as a function of $\cos\theta$ at fixed $k_{1}=k_{2}$. The error bars denote the standard deviation of the mean from 100 different simulation runs.
The second peak at $k\approx 13$ is subtler ($S_{\text{2nd max}}\approx 1.34$ from MC). Rosenfeld again overshoots ($S_{\text{2nd max}}\approx 1.42$), as do the Verlet closure, Malijevský’s functional, WBI, and WBII, in order of increasing accuracy. The MHNC relation reproduces the MC data with very high accuracy, which is perhaps not surprising since it incorporates a high-fidelity MC correlation data parametrization. Beyond the second peak, all curves converge toward $S_{2}(k)\to 1$, with decreasing model-to-model variation.
@@
At low wavenumbers, all FMT formulations provide excellent agreement with MC data. The agreement becomes less precise as $k$ increases. For example, in panels (b–c) and (e–f), where $k_{1}$ spans intermediate to high values ($k_{1}\approx 7-14$), the relative accuracy of the FMT predictions varies drastically with the angle and wavenumbers, showing no clear ‘best’ theory. Interestingly, in the small-angle isosceles configuration with wavenumbers around the main diffraction peak $k_{1}=k_{2}=7.0$, $\cos\theta\approx 1$, all theories show qualitative disagreement with the MC data. We surmise that this is related to earlier observations that the FMTs do not strictly enforce zero probability upon particle overlap [20, 57].
Figure 4: Triplet direct correlation function $\rho^{2}c_{3}(k_{1},k_{2},\theta)$ for a hard-sphere fluid at $\eta=0.49$. Black symbols show Monte Carlo (MC) results. (a) Direct triplet correlations $c_{3}$ against $k$ for $k_{1}=k_{2}=k$ and $\theta=0.25\pi$. (b–d) $c_{3}$ as a function of $\cos\theta$ for various $k_{1}=k_{2}$ values. The error bars denote the standard deviation of the mean from 100 different simulation runs.
Figure 4 presents the corresponding triplet direct correlation function $c_{3}(k_{1},k_{2},\theta)$, plotted as a function of either $k_{1}=k_{2}=k$ or $\cos\theta$. At low wavenumbers (panel a), all FMT functionals describe the MC data well, with the best predictions given by the WBI/II(t) and M theories. This superiority of WBI/I(t) and M is expected, as the low-$k$ regime is primarily governed by thermodynamic consistency and bulk equation-of-state constraints, which these theories incorporate more accurately.
In contrast, at values of $k$ around the main diffraction peak (panel c), discrepancies between the models become more pronounced. In the isosceles configuration ($k=k_{1}=k_{2}=7.0$), the predicted $c_{3}$ curves differ in both amplitude and angular dependence from the MC data. While all functionals capture the overall magnitude of $c_{3}$, only the functionals that incorporate tensorial contributions (T and WBIIt) show reasonable qualitative agreement with the data. This mirrors earlier findings from the pair correlation analysis, where improved thermodynamic functionals gave more accurate $c_{2}(k)$ at small $k$ but not necessarily so at larger $k$. If we increase $k$ to a value of 14.0, all FMTs predict that direct triplet correlations are only relevant for $\cos(\theta)\approx-1$, which is not in agreement with MC data, which show marked oscillations with $\theta$.
@@
Figure 5: Four-body static structure factor $S_{4}(k_{1},k_{2},k_{3},\theta_{12},\theta_{13},\phi_{23})$ for a hard-sphere fluid at $\eta=0.49$. Monte Carlo data (black symbols) are compared to various FMT predictions and the convolution approximation obtained with the Rosenfeld $S_{2}$ (dotted). The wavenumbers are fixed at $k_{1}=k_{2}=7.0$, while $k_{3}\in\{2.0,7.0\}$ as labeled. Each panel corresponds to a different angular configuration. The error bars denote the standard deviation of the mean from 100 different simulation runs.
To assess the predictive power of FMT in capturing higher-order correlations, we now turn to the four-point static structure factor $S_{4}(k_{1},k_{2},k_{3},\theta_{12},\theta_{13},\phi_{23})$. Figure 5 shows comparisons between Monte Carlo (MC) simulations and FMT predictions for six different angular configurations, with $k_{1}=k_{2}=7.0$, and varying $k_{3}\in\{2.0,7.0\}$. The angles $\theta_{12}$ and $\theta_{13}$ are chosen from $\pi/3$ and $2\pi/3$ because these configurations yielded the largest observed correlations. We expect this is due to the appearance of these angles in hexagonal (FCC/HCP), tetrahedral, and icosahedral motifs prevalent in hard-sphere fluids [58, 59, 60, 61].
Across all panels, FMT successfully captures the general shape and location of the principal features of $S_{4}$. However, as with the triplet correlations, different functionals vary significantly in their quantitative accuracy. The original Rosenfeld (R) formulation tends to overpredict the peak heights in nearly all cases [with an exception shown in panel (c)], while White Bear II and tensorial variants consistently under-predict the correlations. Overall (also considering many other configurations not shown here), we have found that the White Bear I functional gives the best results out of all the functionals for the four-body structure. We believe, however, that this is mostly fortuitous.
Interestingly, the convolution approximation (blue dotted lines), which neglects both $c_{3}$ and $c_{4}$, performs surprisingly well in reproducing qualitatively the overall angular structure at this density, as was noted before [12]. In particular, it gives essentially the same qualitative predictions as the different FMTs. This raises the question of which many-body direct correlation functions are essential for an accurate description of $S_{4}$.
Figure 6: Effect of higher-order direct correlation functions on the four-point structure factor $S_{4}$. Each panel shows $S_{4}$ computed from the Rosenfeld functional using the full four-body FMT expression (blue), with $c_{3}$ but not $c_{4}$ retained (green, 3-conv.), and with both $c_{3}=0$ and $c_{4}=0$ (red, 2-conv.). The wavenumbers are fixed at $k_{1}=k_{2}=7.0$, while $k_{3}\in\{2.0,7.0\}$ as labeled.
@@
Remarkably, the results show that the correction to the convolution approximation arises almost entirely from the inclusion of $c_{3}$. The red curves (no $c_{3}$ or $c_{4}$) consistently misrepresent the magnitude of $S_{4}$ compared to the full theory, while the green curves (only $c_{4}=0$) reproduce the full FMT predictions (blue curves) almost exactly. This holds across all tested configurations, and is especially pronounced in panels (c–d), where the omission of $c_{3}$ leads to a dramatic underestimation of the main peak, while the omission of $c_{4}$ leaves the correlations virtually untouched. These results suggest that in the bulk hard-sphere fluid, four-body correlations are dominated by three-body contributions, and that including $c_{4}$ yields only marginal improvements.
@@
To further test the performance of FMT, we consider supercooled liquids, extending our analysis to volume fractions beyond the freezing point. To this end, we simulate a slightly polydisperse quasi-hard-sphere system, as introduced in Ref. [62], which avoids crystallization and allows us to probe the structure of a metastable, amorphous fluid up to $\eta=0.58$. At this high density, corresponding to the supercooled regime, growing structural correlations are known to emerge [4, 63, 64, 65, 66, 67, 68, 69, 70], and four-point structure factors provide a natural way to probe such effects [12].

### v2

@@ Abstract
We study many-body correlation functions within various Fundamental Measure Theory (FMT) formulations and compare their predictions to Monte Carlo simulations of hard-sphere fluids. FMT accurately captures the qualitative behavior of three- and four-body structure, particularly at low and intermediate wavevectors. At higher wavevectors, the predictions of FMT vary in quantitative accuracy. We show that the dominant contributions to the four-point structure factor arise from direct triplet correlations, allowing the evaluation of four-point correlations to be greatly simplified. In glass-forming liquids at high volume fractions, FMT correctly reproduces deviations from the convolution approximation, highlighting FMT’s ability to capture growing structural multipoint correlations upon supercooling.
@@ Introduction
Here $\alpha$ is a label that identifies the particular weight function $\omega_{\alpha}$. The $n$-body direct correlation functions $c_{n}$ are key quantities in density functional theory and FMT, as they describe the response of the system to density perturbations. These functions are derived from the excess free energy functional. In turn, they are related to the $n$-body density through the Ornstein-Zernike relations.
@@
In the homogeneous single-component bulk phase, the two-, three-, and four-body Ornstein-Zernike relations between $c_{n}$ and $S_{n}$ are [42, 12]
@@
$\begin{split}S_{4}&(\mathbf{k}_{1},\mathbf{k}_{2},\mathbf{k}_{3})=\\ &\left(1+\rho^{2}c_{3}(\mathbf{k}_{1},\mathbf{k}_{2})\right)S_{2}(k_{1})S_{2}(k_{2})S_{3}(\mathbf{k}_{1}+\mathbf{k}_{2},\mathbf{k}_{3})\\ +&\left(1+\rho^{2}c_{3}(\mathbf{k}_{1},\mathbf{k}_{3})\right)S_{2}(k_{1})S_{3}(\mathbf{k}_{1}+\mathbf{k}_{3},\mathbf{k}_{2})S_{2}(k_{3})\\ +&\left(1+\rho^{2}c_{3}(\mathbf{k}_{2},\mathbf{k}_{3})\right)S_{3}(\mathbf{k}_{2}+\mathbf{k}_{3},\mathbf{k}_{1})S_{2}(k_{2})S_{2}(k_{3})\\ -&\left(2-\rho^{3}c_{4}(\mathbf{k}_{1},\mathbf{k}_{2},\mathbf{k}_{3})\right)\\ &\quad\times S_{2}(k_{1})S_{2}(|\mathbf{k}_{1}+\mathbf{k}_{2}+\mathbf{k}_{3}|)S_{2}(k_{2})S_{2}(k_{3}),\end{split}$ (10)
where we have explicitly set $\mathbf{k}_{n}=-\sum_{j=1}^{n-1}\mathbf{k}_{j}$ and removed it as a function argument to enforce translational invariance in bulk. These expressions provide the basis for computing three- and four-body correlations and structure factors with FMT. The higher-order structure factors are often approximated purely in terms of $S_{2}$ by assuming that $c_{n>2}=0$, an approach known as the convolution approximation [43].
@@
Kierlik and Rosinberg [44] proposed an alternative formulation that uses only scalar weight functions. In this version, the vector contributions are absorbed into redefined scalar weighted densities so that the same bulk thermodynamics is recovered. Because it recovers exactly the correlation functions generated by Rosenfeld’s original FMT [45], the formulation of Kierlik and Rosinberg is not treated separately here.
@@
$\begin{split}&\Phi_{\text{T}}=-n_{0}\ln(1-n_{3})+\frac{n_{1}n_{2}-\mathbf{n}_{\mathrm{v}1}\cdot\mathbf{n}_{\mathrm{v}2}}{1-n_{3}}\\ &+\frac{n_{2}^{3}-3n_{2}\,\mathbf{n}_{\mathrm{v}2}\cdot\mathbf{n}_{\mathrm{v}2}+\frac{9}{2}\left[\mathbf{n}^{T}_{\mathrm{v}2}\mathbf{n}_{\mathrm{m}2}\mathbf{n}_{\mathrm{v}2}-\mathrm{tr}\left({\mathbf{n}_{\mathrm{m}2}^{3}}\right)\right]}{24\pi(1-n_{3})^{2}}.\end{split}$ (16)
@@
Later, Malijevskỳ [22] (M) proposed a free energy density that reproduces the equation of state of Boublík [50], showing very accurate density profiles near a hard wall. This free energy density reads
@@
Recently, Lutsko [51] has generalized Tarazona’s expression to enforce global stability, in the sense that the free energy must be bounded from below for any density profile. This leads to a functional form given by
$\begin{split}\Phi_{\text{L}}&=-n_{0}\ln(1-n_{3})+\frac{n_{1}n_{2}-\mathbf{n}_{\mathrm{v}1}\cdot\mathbf{n}_{\mathrm{v}2}}{1-n_{3}}\\ &+\frac{\frac{8A+2B}{9}n_{2}^{3}-2An_{2}\,\mathbf{n}_{\mathrm{v}2}\cdot\mathbf{n}_{\mathrm{v}2}+3A\mathbf{n}^{T}_{\mathrm{v}2}\mathbf{n}_{\mathrm{m}2}\mathbf{n}_{\mathrm{v}2}}{24\pi(1-n_{3})^{2}}\\ &+\frac{-(A+B)n_{2}\mathrm{tr}\left(\mathbf{n}_{\mathrm{m}2}^{2}\right)+(2B-A)\mathrm{tr}\left({\mathbf{n}_{\mathrm{m}2}^{3}}\right)}{24\pi(1-n_{3})^{2}},\end{split}$ (18)
which depends on the free parameters $A$ and $B$. Upon the choice (A,B)=(3/2,3/2)(A,B)=(3/2,-3/2), $\Phi_{\text{L}}$ reduces to Tarazona’s expression. Lutsko has shown that this functional is only globally stable if $A,B\geq 0$, and proposes the choice $(A,B)=(1,0)$. In a follow-up work Gül et al. [52] suggests an alternative choice (A,B)=(1.3,1.0)(A,B)=(1.3,-1.0) based on the best incorporation of thermodynamical sum rules. The free energies obtained after substitution of the parameter choices are respectively denoted by (L) and (G).
@@
where $j_{0}(x)=\sin(x)/x$, $j_{1}(x)=(\sin(x)-x\cos(x))/x^{2}$, and $j_{2}(x)=(3\sin(x)-3x\cos(x)-x^{2}\sin(x))/x^{3}$ are the first three spherical Bessel functions of the first kind, $\mathbf{r}\mathbf{r}$ and $\mathbf{k}\mathbf{k}$ are dyadic products, and $\mathbf{I}$ is the unit tensor. We note that by opting to work with the traceless tensorial weight $\omega_{\mathrm{m}2}$, the forms of the tensorial functionals above are slightly different than if they had been written in terms of $\omega_{\mathrm{m}2}^{\prime}=\mathbf{rr}\delta(r-R)/r^{2}$, which is also used in the literature. Additionally, for the vector- and tensor-like weight functions, not only the sign but also the direction of the wave vector is important for the correctness of the result.
The Fourier representations of the weight functions are now used to construct the direct correlation functions in Fourier space using Eq. (6), which can then be used to predict the $n$-body structure factors. This procedure yields fully explicit symbolic expressions, and while they can become very long, their evaluation is essentially free in a computational sense.
@@ Results
To assess the accuracy of various FMT formulations in capturing $n$-body structural correlations, we compute the direct correlation functions $c_{n}(k)$ and static structure factors $S_{n}(k)$ for a hard-sphere fluid in which the particles have diameter $\sigma=1$, which serves as our unit of length. We consider a number density $\rho=0.94$ just below freezing, corresponding to a volume fraction $\eta=\pi\rho/6\approx 0.49$. For reference, we compare each of the theories to correlation functions measured from standard Metropolis Monte Carlo (MC) simulations of $N=30,000$ particles in the canonical ensemble (NVT). To get sufficient statistics, we perform 100 independent equilibration and production runs of $10^{6}$ MC sweeps each, and store a snapshot every $10^{3}$ sweeps during production. Each independent simulation took around 10 CPU hours. From the $10^{5}$ total snapshots, we compute the structure factors $S_{n}$ directly from their definition, Eq. (7) (see Ref. [12] and its supplementary information for more details). In all figures pertaining to the hard-sphere MC simulations, the error bars indicate the standard deviation of the mean across the 100 independent runs. The simulation, analysis, and plotting code can be found in the repository of Ref. [53]. We start with a comparison for $n=2$, and consider n=3, 4n=3,\ 4 later in this section.
@@
Figure 1: (a) Fourier-space two-body direct correlation function $\rho c_{2}(k)$ for a hard-sphere fluid at volume fraction $\eta=0.49$, comparing various theoretical models to Monte Carlo (MC) simulation data (black symbols). (b) Magnification of the low-$k$ region. FMT-based predictions are shown for the original Rosenfeld (R), White Bear I (WBI), White Bear II (WBII), Malijevský (M) Lutsko (L) and Gül (G) functionals, alongside results obtained from Ornstein–Zernike closures, specifically the Verlet (V) and Modified Hypernetted Chain (MHNC) theories. For this and all following figures, the error bars denote the standard deviation of the mean from 100 different simulation runs. Here, the error bars are typically smaller than the markers.
To place the performance of the different FMT formulations for the 2-point bulk structure in theoretical context, we compare the FMT predictions here additionally to the very successful Verlet (V) [54, 55] and Modified Hypernetted Chain (MHNC) closures of the Ornstein-Zernike relation [56], which are go-to methods for predicting pair structure in bulk. The MHNC closure uses a parametrization of the bridge function of hard spheres measured from Monte Carlo (MC) simulations and is therefore expected to offer virtually exact predictions for $r>1$ [57, 58]. The integral equation theories are solved by the software introduced in Ref. [58].
@@
Figure 1 shows the two‐body direct correlation function $\rho c_{2}(k)$ for a hard‐sphere fluid at volume fraction $\eta=0.49$ comparing Monte Carlo data to a range of theoretical models: Rosenfeld’s original fundamental‐measure theory (R), White Bear I (WBI), White Bear II (WBII), Malijevský’s variant (M), the newer functionals of Lutsko (L) and Gül (G), and the two integral‐equation closures. The tensorial theories (T, WBIIt) are not considered here because their tensorial contributions have no effect at the pair level, reducing their predictions to those of their non-tensorial equivalents (R, WBII). Qualitatively, all FMTs show the correct behavior for the direct correlation function. Quantitatively, for the longest wavelengths ($k\to 0$), Rosenfeld’s FMT predicts a considerably more negative $c_{2}(k)$ than the others. This highlights its commonality with Percus–Yevick theory, resulting in an isothermal compressibility that is lower than that computed from MC data [1]. The FMTs of Lutsko and Gül perform better, but still show significant errors in the low-$k$ regime. In contrast, the White Bear I/II functionals, which enforce the Boublík–Mansoori–Carnahan–Starling–Leland or generalized Carnahan–Starling equation of state [46, 47, 48], bring $c_{2}(0)$ into near‐perfect agreement with simulation. Malijevský’s free energy density, also built on a high‐accuracy bulk equation of state, gives essentially the same low‐$k$ result as WBII, very slightly underestimating $c_{2}(0)$ compared to WBI/II and the MC data. The Verlet closure, likewise, matches the MC compressibility almost exactly, whereas the MHNC closure fails to accurately capture the small-$k$ behavior because its parametrization was performed for $r>1$ only. At high $k$ ($k\gtrsim 6$), every model’s $c_{2}(k)$ approaches zero in lockstep with MC. To visualize the differences between the theories, we consider the static structure factor $S(k)=1/(1-\rho c_{2}(k))$, which amplifies the small differences in $\rho c(k)$ at larger $k$.
Figure 2 presents the static structure factor $S_{2}(k)$. The principal diffraction peak at $k\approx 7$ attains $S_{\text{max}}\approx 2.9$ in simulation. Rosenfeld, Lutsko, and Gül overestimate this peak ($S_{\text{max}}\approx 3.1$, $S_{\text{max}}\approx 3.5$, and , $S_{\text{max}}\approx 3.0$, resp.), while WBI and WBII underestimate it ($S_{\text{max}}\approx 2.7$ and $S_{\text{max}}\approx 2.4$, resp.). Malijevský’s form produces much better agreement with simulation data. The Verlet closure matches the first peak with virtually no error, and so does MHNC.
Figure 3: Triplet structure factor $S_{3}(k_{1},k_{2},\theta)$ at $\eta=0.49$ from Monte Carlo (MC) simulations (black circles) compared to predictions from the different FMTs. The convolution approximation ($c_{3}=0$) is also shown, for which we have used the Rosenfeld prediction of $S_{2}$. Top row: $S_{3}$ as a function of $k_{1}$ at fixed $\theta$ and $k_{2}$. Bottom row: $S_{3}$ as a function of $\cos\theta$ at fixed $k_{1}=k_{2}$.
The second peak at $k\approx 13$ is subtler ($S_{\text{2nd max}}\approx 1.34$ from MC). Rosenfeld again overshoots ($S_{\text{2nd max}}\approx 1.42$), as do the Verlet closure, Malijevský’s functional, WBI, G, and WBII, in order of increasing accuracy. Lutsko’s FMT slightly undershoots, and the MHNC relation reproduces the MC data with very high accuracy, which is perhaps not surprising since it incorporates a high-fidelity MC correlation data parameterization. Beyond the second peak, all curves converge toward $S_{2}(k)\to 1$, with decreasing model-to-model variation.
@@
At low wavenumbers, all FMT formulations provide excellent agreement with MC data. The agreement becomes less precise as $k$ increases. For example, in panels (b–c) and (e–f), where $k_{1}$ spans intermediate to high values ($k_{1}\approx 7-14$), the relative accuracy of the FMT predictions varies drastically with the angle and wavenumbers, showing no clear ‘best’ theory. Interestingly, in the small-angle isosceles configuration with wavenumbers around the main diffraction peak $k_{1}=k_{2}=7.0$, $\cos\theta\approx 1$, all theories show qualitative disagreement with the MC data. We surmise that this is related to earlier observations that the FMTs do not strictly enforce zero probability upon particle overlap [20, 61]. See also the work of Coslovich [42] for similar observations.
Figure 4: Triplet direct correlation function $\rho^{2}c_{3}(k_{1},k_{2},\theta)$ for a hard-sphere fluid at $\eta=0.49$. Black symbols show Monte Carlo (MC) results. (a) Direct triplet correlations $c_{3}$ against $k$ for $k_{1}=k_{2}=k$ and $\theta=0.25\pi$. (b–d) $c_{3}$ as a function of $\cos\theta$ for various $k_{1}=k_{2}$ values.
Figure 4 presents the corresponding triplet direct correlation function $c_{3}(k_{1},k_{2},\theta)$, plotted as a function of either $k_{1}=k_{2}=k$ or $\cos\theta$. At low wavenumbers (panels a, b), all FMT functionals describe the MC data well, with the best predictions given by the WBI/II(t) and M theories. This superiority of WBI/I(t) and M is expected, as the low-$k$ regime is primarily governed by thermodynamic consistency and bulk equation-of-state constraints, which these theories incorporate more accurately.
In contrast, at values of $k$ around the main diffraction peak (panel c), discrepancies between the models become more pronounced. In the isosceles configuration ($k=k_{1}=k_{2}=7.0$), the predicted $c_{3}$ curves differ in both amplitude and angular dependence from the MC data. While all functionals capture the overall magnitude of $c_{3}$, only the functionals that incorporate tensorial contributions (T, WBIIt, L, and G) show reasonable qualitative agreement with the data. This mirrors the findings from our pair level analysis above, where we have shown that improved thermodynamic functionals give more accurate $c_{2}(k)$ at small $k$ but not necessarily so at larger $k$. If we increase $k$ to a value of 14.0, all FMTs predict that direct triplet correlations are only relevant for $\cos(\theta)\approx-1$, which is not in agreement with MC data, which show marked oscillations with $\theta$.
@@
Figure 5: Four-body static structure factor $S_{4}(k_{1},k_{2},k_{3},\theta_{12},\theta_{13},\phi_{23})$ for a hard-sphere fluid at $\eta=0.49$. Monte Carlo data (black symbols) are compared to various FMT predictions and the convolution approximation obtained with the Rosenfeld $S_{2}$ (black, dashed). The wavenumbers are fixed at $k_{1}=k_{2}=7.0$, while $k_{3}\in\{2.0,7.0\}$ as labeled. Each panel corresponds to a different angular configuration.
To assess the predictive power of FMT in capturing higher-order correlations, we now turn to the four-point static structure factor $S_{4}(k_{1},k_{2},k_{3},\theta_{12},\theta_{13},\phi_{23})$. Figure 5 shows comparisons between Monte Carlo (MC) simulations and FMT predictions for six different angular configurations, with $k_{1}=k_{2}=7.0$, and varying $k_{3}\in\{2.0,7.0\}$. The angles $\theta_{12}$ and $\theta_{13}$ are chosen from $\pi/3$ and $2\pi/3$ because these configurations yielded the largest observed correlations. We expect this is due to the appearance of these angles in hexagonal (FCC/HCP), tetrahedral, and icosahedral motifs prevalent in hard-sphere fluids [62, 63, 64, 65]. Apart from the six wave vector configurations shown here, we have tested another 92 different sets of angles and wave numbers. The corresponding figures are available at Ref. [53].
Across all panels, FMT successfully captures the general shape and location of the principal features of $S_{4}$ with minor exceptions. However, as with the triplet correlations, different functionals vary significantly in their quantitative accuracy. The original Rosenfeld (R) formulation and Lutsko’s FMT tend to overpredict the peak heights in nearly all cases, while White Bear II and its tensorial version consistently underpredict the correlations. Overall (also considering many other configurations not shown here), we have found that the White Bear I functional gives the best results out of all the functionals for the four-body structure. We believe, however, that this is mostly fortuitous.
Interestingly, the convolution approximation (black dashed lines), which neglects both $c_{3}$ and $c_{4}$, performs surprisingly well in reproducing qualitatively the overall angular structure at this density, as was noted before [12]. In particular, it gives essentially the same qualitative predictions as the different FMTs. This raises the question of which many-body direct correlation functions are essential for an accurate description of $S_{4}$.
Figure 6: Effect of higher-order direct correlation functions on the four-point structure factor $S_{4}$. Each panel shows $S_{4}$ computed from the Rosenfeld functional using the full four-body FMT expression (solid lines), with $c_{3}$ but not $c_{4}$ retained (dotted, 3-conv.), and with both $c_{3}=0$ and $c_{4}=0$ (dashed, 2-conv.). The wavenumbers are fixed at $k_{1}=k_{2}=7.0$, while $k_{3}\in\{2.0,7.0\}$ as labeled.
@@
Remarkably, the results show that the correction to the convolution approximation arises almost entirely from the inclusion of $c_{3}$. The dashed curves (no $c_{3}$ or $c_{4}$) consistently misrepresent the magnitude of $S_{4}$ compared to the full theory, while the dotted curves (only $c_{4}=0$) reproduce the full FMT predictions (green solid curves) almost exactly. This holds across all tested configurations, and is especially pronounced in panels (c–d), where the omission of $c_{3}$ leads to a dramatic underestimation of the main peak, while the omission of $c_{4}$ leaves the correlations virtually untouched. These results suggest that in the bulk hard-sphere fluid, four-body correlations are dominated by three-body contributions, and that including $c_{4}$ yields only marginal improvements.
@@
To further test the performance of FMT, we consider supercooled liquids, extending our analysis to volume fractions beyond the freezing point. To this end, we simulate a 6% polydisperse quasi-hard-sphere system, as introduced in Ref. [66], which avoids crystallization and allows us to probe the structure of a metastable, amorphous fluid up to volume fractions of $\eta=0.58$. In this system, the $N=1000$ particles interact with an inverse power-law potential $u(r)\propto r^{-36}$. At this high density, corresponding to the supercooled regime, growing structural correlations are known to emerge [4, 67, 68, 69, 70, 71, 72, 73, 74], and four-point structure factors provide a natural way to probe such effects [12].